# Transfer of Knowledge through Reverse Annealing: A Preliminary Analysis of the Benefits and What to Share

### v1

@@ 3 Method
RA is a specific type of quantum annealing designed to perform local refinement of good states found elsewhere (15). To achieve this, RA resorts to a time-dependent Hamiltonian:
@@
Therefore, the RA Schedule can be defined as (18):
@@ 4 Experimentation & Discussion
A fixed RA schedule has been used for all the tests, [(0.0, 1.0), (2.5, 0.5), (102.5, 0.5), (102.75, 1.0)], which has been obtained through an empirical procedure performed in lab. For the sake of fair comparison, the annealing_time of all the forward annealing processes has been adjusted to the RA schedule duration so that all runs carried out in this research access the quantum computer the same amount of time. We recommend papers such as (18) to readers interested in the analysis of the reverse annealing schedule. Lastly, reinitialize_state=True.
For each (source instance - target instance) combination, 10 independent runs have been executed, with Table 2 depicting the best result found among these executions, along with the average and the standard deviation. As an example, 10 independent runs have been executed to solve s14, using the best solution found for s14_0.2_L2L as the input bitstring in the RA process (i.e., source of knowledge). For this combination, the energy of the best solution found is -12421, while the average and standard deviation are -12420.2 and 0.74, respectively. For comparison purposes, we also depict in Table 2 the baseline results obtained by s14 and s16 by means of forward annealing. In these cases, the source of knowledge has been represented as “–”.
Table 2: Results of the Reverse Annealing experimentation. Optimal results, obtained using the industry-oriented Quantagonia’s Hybrid Solver (6): s14=12422; s16=10718. In bold results that equalize or improve the performance of the forward annealing baseline (which is -12422 for s14 and -10716 for s16).
@@
The results shown in this table provide answers to RQ1. If we pay attention to s14, we could preliminarily conclude that the sharing of knowledge does not imply an improvement with respect to the best solution found by the annealer. This reflection cannot be replicated if we look at the values relative to the mean and standard deviation, where the use of solutions from descendant-instances clearly implies better performance. In summary, the sharing of knowledge through RA leads to more reliable and robust behavior of the annealer in the case of s14.
@@ 5 Conclusions & Further Work
  • Conduct a more in-depth study, including a comparison of computational resources and time and using other optimization problems with benchmarking purposes, such as the Traveling Salesman Problem or the Bin Packing Problem. Potential limitations when applying RA to more real-world scenarios will also be studied.
@@
  • Finally, inspired by the findings described in (36), further experiments will be carried out with larger instances, with the intention of determining whether the conclusions presented in this study are replicable in such cases.

### v2

@@ 3 Method
As mentioned in the introduction, a QA is a specialized quantum device engineered to tackle optimization problems through a process inspired by classical Simulated Annealing. More specifically, a QA operates on the principle of adiabatic computation, where an initial, easily prepared Hamiltonian is gradually evolved from its ground state to the ground state of a final, problem-specific Hamiltonian. If this evolution is sufficiently slow, the adiabatic theorem ensures that the system stays in the ground state throughout the computation. In QA, the adiabatic theorem is deliberately relaxed, permiting the process to evolve more quickly than the adiabatic limit would normally permit. Consequently, transitions to higher energy states often occur during the evolution. Despite this computational model can also be considered also universal (34), the D-Wave annealer, which is the one used in this study, is based on an Ising Hamiltonian, which limits the kind of problems that can be executed on the computer. Nevertheless, this type of QA is well-suited for solving combinatorial optimization problems as the one addressed in this paper (35).
Regarding the RA method, it is a specific type of quantum annealing process designed to perform local refinement of good states found elsewhere (15). To achieve this, RA resorts to a time-dependent Hamiltonian:
@@
Therefore, the RA Schedule can be defined as shown in the following Equation 2 (18):
@@ 4 Experimentation & Discussion
A fixed RA schedule has been used for all the tests, [(0.0, 1.0), (2.5, 0.5), (102.5, 0.5), (102.75, 1.0)], which has been obtained through an empirical procedure performed in lab. More precisely, an experimentation was carried out in which 25 different RA schedule were tested. Minor adjustments were made to the $s_{p}$, $t_{r}$, $t_{p}$, and $t_{q}$ parameters. To assess the quality of the designed RA schedule, each one was used to solve the two target-instance 10 times each, using one random solution for the problem as input. The aforementioned schedule was selected due to its superior performance compared to the others.
For the sake of fair comparison, the annealing_time of all the forward annealing processes has been adjusted to the RA schedule duration so that all runs carried out in this research access the quantum computer the same amount of time. We recommend papers such as (18) to readers interested in the analysis of the reverse annealing schedule. Lastly, reinitialize_state=True.
For each (source instance - target instance) combination, 10 independent runs have been executed, with Table 2 depicting the best result found among these executions, along with the average and the standard deviation. As an example, 10 independent runs have been executed to solve s14, using the best solution found for s14_0.2_L2L as the input bitstring in the RA process (i.e., source of knowledge). For this combination, the energy of the best solution found is -12421, while the average and standard deviation are -12420.2 and 0.74, respectively. For comparison purposes, we also depict in Table 2 the baseline results obtained by s14 and s16 by means of forward annealing. In these cases, the source of knowledge has been represented as “–”. It is important to highlight that the optimal results for each instance have been achieved using the industry-oriented Quantagonia Hybrid Solver (QHS, (6)). In essence, this algorithm is divided into two distinguishable steps: Initially, it executes a set of primal heuristics to solve the entire problem. These heuristics can be classical or quantum, utilizing external services. Subsequently, QHS enhances the best solution identified by the primal heuristics through a classical Branch-and-Bound algorithm. One of the key strengths of QHS is the inclusion of an optimality proof mechanism, which provides the optimality gap along with the results. The optimality gap indicates the remaining potential in the optimization process. This feature is the primary reason we selected QHS as the solver to achieve optimal results for the considered instances.
Table 2: Results of the Reverse Annealing experimentation. Optimal results, obtained using the industry-oriented Quantagonia’s Hybrid Solver: s14=12422; s16=10718. In bold results that equalize or improve the performance of the forward annealing baseline (which is -12422 for s14 and -10716 for s16).
@@
The results shown in Table 2 provide answers to RQ1. If we pay attention to s14, we could preliminarily conclude that the sharing of knowledge does not imply an improvement with respect to the best solution found by the annealer. This reflection cannot be replicated if we look at the values relative to the mean and standard deviation, where the use of solutions from descendant-instances clearly implies better performance. In summary, the sharing of knowledge through RA leads to more reliable and robust behavior of the annealer in the case of s14.
@@ 5 Conclusions & Further Work
  • $\blacksquare$
    Future work of considerable importance will likely focus on advancing towards the generalization of the conclusions drawn from this study. To reach this ambitious goal, the following steps are essential:
    • Inspired by the findings described in (38), further experiments should be carried out with larger instances, with the intention of determining whether the conclusions are replicable in such cases.
    • Further study the impact of the RA Schedules on knowledge transfer. To do this, the preliminary study already described in Section 4 will be revisited, and the knowledge transfer paradigm will be examined under the mentioned 25 RA Schedules.
    • Carry out a thorough analysis of computational resources and time.
    • Examine whether the findings from the experimentation can be replicated in other optimization problems, such as the Traveling Salesman Problem or the Bin Packing Problem.
    • Study the potential limitations when applying RA to more real-world scenarios.
  • $\blacksquare$
    Fairly compare RA as a local refinement method (using the same instance for source- and target- tasks) with RA as a transfer knowledge procedure.
  • $\blacksquare$